X - ray spectrum of a cobalt largest ( Z = 27 ) contains strong K(Alpha) line of wavelength
0.1785 nm and a weak line having wavelength of 0.1537 nm due to impurities. The atomic number of impurity element is?
1
Expert's answer
2020-06-16T09:36:29-0400
"\frac{1}{\lambda}=\frac{k(z-1)^2}{1}"

"\frac{\lambda_A}{\lambda_B}=\frac{(z_B-1)^2}{(z_A-1)^2}"

"\frac{0.1785}{0.1537}=\frac{(z_B-1)^2}{(27-1)^2}"

"z_B=29"